# Merlin: A Vision Language Foundation Model for 3D Computed Tomography

### v1

@@ Abstract
Over 85 million computed tomography (CT) scans are performed annually in the US, of which approximately one quarter focus on the abdomen. Given the current shortage of both general and specialized radiologists, there is a large impetus to use artificial intelligence to alleviate the burden of interpreting these complex imaging studies while simultaneously using the images to extract novel physiological insights. Prior state-of-the-art approaches for automated medical image interpretation leverage vision language models (VLMs). However, current medical VLMs are generally limited to 2D images and short reports, and do not leverage electronic health record (EHR) data for supervision. To overcome these shortcomings for abdominal CT interpretation, we introduce Merlin - a 3D VLM that leverages both structured EHR and unstructured radiology reports for supervision without requiring additional manual annotations. We train Merlin using a high-quality clinical dataset of paired CT scans (6+ million images from 15,331 CTs), EHR diagnosis codes (1.8+ million codes), and radiology reports (6+ million tokens). We comprehensively evaluate Merlin on 6 task types and 752 individual tasks. The non-adapted (off-the-shelf) tasks include zero-shot findings classification (31 findings), phenotype classification (692 phenotypes), and zero-shot cross-modal retrieval (image to findings and image to impressions), while model adapted tasks include 5-year chronic disease prediction (6 diseases), radiology report generation, and 3D semantic segmentation (20 organs). We perform internal validation on a test set of 5,137 CTs, and external validation on 7,000 clinical CTs and on two public CT datasets (VerSe, TotalSegmentator). Beyond these clinically-relevant evaluations, we assess the efficacy of various network architectures and training strategies to depict that Merlin has favorable performance to existing task-specific baselines. We derive data scaling laws to empirically assess training data needs for requisite downstream task performance. Furthermore, unlike conventional VLMs that require hundreds of GPUs for training, we perform all training on a single GPU. This computationally efficient design can help democratize foundation model training, especially for health systems with compute constraints. We plan to release our trained models, code, and dataset, pending manual removal of all protected health information.
@@ 1 Main
Over 85 million computed tomography (CT) scans are performed per year in the US for Economic Co-operation and Development (2022); Health (2021); Winder et al. (2021), with approximately 300 million CTs performed globally Schöckel et al. (2020). Amongst these studies, CTs of the abdomen represent approximately one fourth of all examinations performed Kanal et al. (2017). These abdominal CT scans can consist of upwards of 300 slices per series with numerous anatomical structures that need to be examined, leading to time-consuming interpretation, often requiring 20 minutes per exam Udare et al. (2022). Moreover, recent literature suggests that abdominal CT scans contain biomarkers of early diseases that routinely go unreported Bellolio et al. (2017); Liu et al. (2022a); Lee et al. (2022); Thibault et al. (2012); Bates and Pickhardt (2022); Kuriyan (2018); Blankemeier et al. (2022); Zambrano Chaves et al. (2023). With the current volume of medical imaging and a 6% annual News (2024) increase in medical imaging utilization, the burden on radiologists is significant. Nonetheless, over the past decades, the number of radiology residency positions in the US has remained relatively constant at 1011 in 2006 and 1113 in 2020 Wadhwa et al. (2022). This is despite 1,800+ open radiologist positions on the American College of Radiology job board of Radiology (2024) from a pool of roughly 32,000 total active radiologists in the US of Labor Statistics (2024)***Number of job openings on the ACR job board based on accessing the site on May 26, 2024.. With the supply of new radiologists remaining relatively constant and an ever increasing utilization of medical imaging, the radiologist shortage is projected to expand to over 19,000 by 2036 Hudnall (2024); Milburn M (2024); Patel et al. (2020); Rimmer (2017).
Machine learning (ML) has shown promise in various medical imaging tasks Ardila et al. (2019); Eng et al. (2021); Cao et al. (2023); Wang et al. (2024), inspiring optimism about its potential to offset the increasing burden faced by radiologists Hudnall (2024); Langlotz (2023). As of May 2024, of the 882 FDA-cleared ML-enabled devices, 671 (76%) relate to radiology Food and Administration (2024). Despite this large prevalence, the current status-quo of training ML algorithms entails using unimodal (imaging-only) algorithms and retraining for new tasks from scratch using supervised ML with manually curated labels, even if it may be for the same modality or anatomy. In the medical imaging scenario, generating such labels requires expensive clinical expert time, limiting the development of capable AI models for a wide array of tasks.
@@
CLIP-based methods could be readily applied in the clinical setting by training with medical images and corresponding radiology reports that are generated routinely during clinical care, adding no additional data labeling cost. Many institutions, including ours, de-identify this data to maintain patient privacy while providing a unique setting for research on this high-quality human-annotated data. Recent ethical viewpoints describe how large-scale already-acquired clinical data could be responsibly used for secondary purposes such as training ML algorithms, to ensure that the data are used for the benefit of future patients and society Larson et al. (2020). Following such frameworks, VLMs like MedCLIP Wang et al. (2022), BiomedCLIP Zhang et al. (2023a), LLaVA-Med Li et al. (2023a), Med-Flamingo Moor et al. (2023), Med-PaLM M Tu et al. (2023), RadFM Wu et al. (2023), XrayGPT Thawkar et al. (2023), ELIXR Xu et al. (2023), RoentGen Chambon et al. (2022a, b), CheXagent Chen et al. (2024), and MAIRA-1 Hyland et al. (2023) demonstrate the potential of VLMs applied within the radiology domain.
@@
    We develop a training strategy and foundation model called MerlinNamed after the legendary figure said to be able to perceive everything in the present, our model Merlin simultaneously processes the entirety of 3D information in a CT volume. that leverages the structured and unstructured data within hospitals to train an abdominal CT visual model. Using this training strategy, we train an extremely compute-efficient vision model on a single GPU using paired CTs (6,387,231 images from 15,331 CTs), EHR diagnosis codes (1,839,559 codes), and radiology reports (6,036,645 tokens). Merlin allows processing the entire 3D voxel data in a CT image at once.
@@
    We evaluate Merlin on a comprehensive set of 6 task types and 752 individual tasks. The non-adapted (off-the-shelf tasks) include zero-shot findings classification (31 findings), phenotype classification (692 phenotypes), and zero-shot cross-modal retrieval (image to findings and image to impressions), while model adapted tasks include 5-year disease prediction (6 diseases), radiology report generation, and 3D semantic segmentation (20 organs).
@@
    We perform internal validation on 5,137 CTs as well as external validation on 7,000 CTs from another institution, and on two publicly available datasets focused on abdominal CT (VerSe Liebl et al. (2021) and TotalSegmentator Wasserthal et al. (2023)). We demonstrate that our singular Merlin model outperforms carefully chosen task-specific baselines on our suite of benchmarking tasks.
@@
  5. 5.
    We plan to release our trained models, code, and dataset, pending manual review and removal of all personal health information.
@@ 2 Results
We present results across 6 evaluation task types comprising 752 individual tasks, that can be performed of-the-shelf without adaptation (Figure 1b - Figure 1d), or with additional adaptation (Figure 1e - Figure 1g). The non-adapted tasks that we evaluate Merlin on include zero-shot classification (31 classification tasks; Figure 1b), phenotype classification (692 classification tasks; Figure 1c), and zero-shot cross-modal retrieval (image to findings and image to impressions; Figure 1d). The adapted tasks that we evaluate Merlin on include 5-year disease prediction (6 classification tasks; Figure 1e), radiology report generation (Figure 1f), and 3D segmentation (20 organs; Figure 1g).
Figure 1: Overview of Merlin training and evaluation. (a) Merlin training strategy. Diagnosis codes from the EHR are used as labels for Merlin training, with a binary cross entropy loss. Radiology reports are also used for training, with an InfoNCE loss Oord et al. (2018). Training with diagnosis codes and radiology reports is either staged or performed in a multi-task manner. Merlin is then evaluated on non-adapted tasks that can be performed without any architectural or weight modifications. These include (b) zero-shot findings classification, (c) phenotype classification, and (d) zero-shot cross-modal retrieval. Adapting Merlin enables us to perform (e) 5-year disease prediction, (f) radiology report generation, and (g) 3D semantic segmentation. All error bars are 95% confidence intervals.
@@ 2.1 Zero-shot Findings Classification
Zero-shot findings classification assesses Merlin’s ability to classify the presence of common imaging findings based on text prompts that a user can develop, which are likely distinct from prompts seen during training. In Figure 2b we evaluate Merlin zero-shot classification across 30 abdominal CT findings on our internal and external clinical datasets. Merlin achieves an average F1 score of 0.741 (95%CI [0.727-0.755]) on the internal validation dataset and an average F1 score of 0.647 (95%CI [0.607-0.678]) on the external validation dataset, significantly outperforming OpenCLIP Cherti et al. (2022) and BioMedCLIP Zhang et al. (2023b) (p < 0.001) in both settings (Figure 2b). Qualitatively, we observe in Figure 2c that Merlin’s external performance remains high on diseases with coarse-grained or salient features, e.g. pleural effusion, splenomegaly, ascites, surgically absent gallbladder, prostatomegaly, anasarca, and abdominal aortic aneurysm. Performance expectedly decreases on more challenging findings that require more subtle and fine-grained features, e.g. appendicitis, metastatic disease, lymphadenopathy, and free air. We also plot Merlin performance without radiology report splitting (W/O splitting in Figure 2c). Radiology report splitting refers to splitting the radiology reports into sections describing different anatomical structures for subsequent contrastive learning (e.g. "liver: normal" and "vasculature: patent"). Without radiology report splitting, Merlin achieves an average F1 score of 0.656 (95%CI [0.640, 0.671]), similar to the Merlin external evaluation performance. On a separate external dataset (VerSe Löffler et al. (2020) vertebral fracture detection), Merlin achieves a zero-shot F1 of 0.767 (95%CI [0.623-0.867]). In Figure 2d, we establish quantitative relationships to assess how expanding the pretraining dataset would improve zero-shot classification. This analysis helps determine the extent of pretraining data necessary for obtaining a specified zero-shot classification performance. The resulting power law is $\overline{F1}=0.458D^{0.0524}$, with $D$ representing the number of paired pretraining CT imaging and report data.
@@
Figure 2: Zero-shot findings classification. (a) Depicts how zero-shot classification is performed where text embeddings from disease presence prompts and disease absence prompts are compare to the image embedding. (b) We compare the performance of OpenCLIP Cherti et al. (2022), BioMedCLIP Zhang et al. (2023b), Merlin on an internal dataset, and Merlin without radiology report splitting. We further evaluate Merlin on an external clinical dataset and the VerSe Löffler et al. (2020) external fracture detection dataset. (c) Performance of BioMedCLIP, Merlin, and Merlin without report splitting on the internal dataset, as well as Merlin on the external dataset, across 30 findings assessed on abdominal CT scans. (d) Merlin zero-shot classification performance improves with increasing pretraining dataset size. (e) An ablation study across various aspects of Merlin’s pretraining strategy. "Rpt." is shorthand for "report" and indicated training with radiology reports only. Staged (Stg.) refers to performing weakly supervised training with EHR in a first training stage and then training with radiology reports in a second stage. This is in contrast to multi-task learning (MTL) where EHR and radiology reports are used for training simultaneously.
@@ 2.2 Phenotype Classification
In Figure 3b we compute data scaling law curves to assess how Merlin performance improves with respect to training data. Similar to zero-shot classification, we find that increasing training data improves performance. Our analysis also reveals power laws with the following slopes and intercepts: ($\overline{AUROC}=0.479D^{0.0568}$ and $\overline{AUPRC}=0.0157x^{0.239}$).
@@
In Figure 3e, we apply the latent shift counterfactual method Cohen et al. (2021a) to qualitatively investigate the image features that Merlin uses to perform image classification. We examine an instance of “pleural effusion” (left) where the effusion is localized to the left lung and is reduced in the counterfactual, indicating that Merlin is leveraging the features we expect. We also examine “splenomegaly” classification (right). We observe that the size of the spleen is reduced in the counterfactual relative to the original image, adding credence to the validity of imaging features that Merlin uses for image classification.
Figure 3: Phenotype classification. (a) Average AUROC performance for the top 20 phenotype groups listed in order of prevalence (black line). (b) Data scaling law experiments that measure how average AUROC (top) and average AUPRC (bottom) across the 692 phenotypes scale as the amount of pretraining data varies. (c) Average AUROC (left chart) and AUPRC (right chart) across all 692 phenotypes, the top quartile of 173 phenotypes, and the botton quartile of 173 phenotypes across several baseline models. All baseline models are trained using the phenotypes in the pretraining dataset. The dashed lines denote random chance performance. Note that Merlin, which uses the best performing backbone of ResNet152, is further trained using radiology reports. (d) Average AUROC as a function of model stem hyper-parameters. We find that a smaller receptive field yields better performance. (e) Counterfactual analyses of pleural effusion classification (left; image from TCIA Clark et al. (2013)) and splenomegaly classification (image from our internal test set). We annotate the zoomed in images by outlining the pathologies. The red lines border pathologies in the original images. The blue lines border pathologies in the counterfactual images. Counterfactual outlines are drawn over the original images with dotted lines and the original image outlines are also drawn over the counterfactual images with dotted lines. This allows comparing the size and shape of the pathologies between the original images and the counterfactuals, indicating that Merlin is indeed using appropriate features for image classification.

### v2

# Merlin: A Computed Tomography Vision-Language Foundation Model and Dataset

@@ Abstract
Over 85 million computed tomography (CT) scans are performed annually in the US, of which approximately one quarter focus on the abdomen. Given the current shortage of both general and specialized radiologists, there is a large impetus to use artificial intelligence to alleviate the burden of interpreting these complex imaging studies while simultaneously using the images to extract novel physiological insights. Prior state-of-the-art approaches for automated medical image interpretation leverage vision language models (VLMs) that utilize both the image and the corresponding textual radiology reports. However, current medical VLMs are generally limited to 2D images and short reports. To overcome these shortcomings for abdominal CT interpretation, we introduce Merlin - a 3D VLM that leverages both structured electronic health records (EHR) and unstructured radiology reports for pretraining without requiring additional manual annotations. We train Merlin using a high-quality clinical dataset of paired CT scans (6+ million images from 15,331 CTs), EHR diagnosis codes (1.8+ million codes), and radiology reports (6+ million tokens) for training. We comprehensively evaluate Merlin on 6 task types and 752 individual tasks, covering a variety of diagnostic, prognostic, and quality-related tasks. The non-adapted (off-the-shelf) tasks include zero-shot findings classification (31 findings), phenotype classification (692 phenotypes), and zero-shot cross-modal retrieval (image to findings and image to impressions), while model adapted tasks include 5-year chronic disease prediction (6 diseases), radiology report generation, and 3D semantic segmentation (20 organs). We perform internal validation on a test set of 5,137 CTs, and external validation on 44,098 CTs from three external sites spanning both abdomen and chest anatomies, and on two public CT datasets (VerSe, TotalSegmentator). Beyond these clinically-relevant evaluations, we assess the efficacy of various network architectures and training strategies to depict that Merlin has favorable performance to existing task-specific baselines. We derive data scaling laws to empirically assess training data needs for requisite downstream task performance. Our resource-friendly approach enables health systems to train their own foundation models with modest computational resources. We also release our trained models, code, and dataset, available at: https://github.com/StanfordMIMI/Merlin.
@@ 1 Main
Over 85 million computed tomography (CT) scans are performed per year in the US for Economic Co-operation and Development ; Health (2021); Winder et al. (2021), with approximately 300 million CTs performed globally Schöckel et al. (2020). Amongst these studies, CTs of the abdomen represent approximately one fourth of all examinations performed Kanal et al. (2017). These abdominal CT scans can consist of upwards of 300 slices per series with numerous anatomical structures that need to be examined, leading to time-consuming interpretation, often requiring 20 minutes per exam Udare et al. (2022). Moreover, recent literature suggests that abdominal CT scans contain biomarkers of early diseases that routinely go unreported Bellolio et al. (2017); Liu et al. (2022a); Lee et al. (2022); Thibault et al. (2012); Bates and Pickhardt (2022); Kuriyan (2018); Blankemeier et al. (2022); Zambrano Chaves et al. (2023). With the current volume of medical imaging and a 6% annual News (2024) increase in medical imaging utilization, the burden on radiologists is significant. Nonetheless, over the past decades, the number of radiology residency positions in the US has remained relatively constant at 1011 in 2006 and 1113 in 2020 Wadhwa et al. (2022). This is despite 1,800+ open radiologist positions on the American College of Radiology job board of Radiology (2024) from a pool of roughly 32,000 total active radiologists in the US of Labor Statistics (2024)***Number of job openings on the ACR job board based on accessing the site on May 26, 2024.. With the supply of new radiologists remaining relatively constant and an ever increasing utilization of medical imaging, the radiologist shortage is projected to expand to over 19,000 by 2036 Hudnall ; Milburn M ; Patel et al. (2020); Rimmer (2017).
Machine learning (ML) has shown promise in various medical imaging tasks Ardila et al. (2019); Eng et al. (2021); Cao et al. (2023); Wang et al. (2024), inspiring optimism about its potential to offset the increasing burden faced by radiologists Hudnall ; Langlotz (2023). As of May 2024, of the 882 FDA-cleared ML-enabled devices, 671 (76%) relate to radiology Food and Administration . Despite this large prevalence, the current status-quo of training ML algorithms entails using unimodal (imaging-only) algorithms and retraining for new tasks from scratch using supervised ML with manually curated labels, even if it may be for the same modality or anatomy. In the medical imaging scenario, generating such labels requires expensive clinical expert time, limiting the development of capable AI models for a wide array of tasks.
@@
CLIP-based methods could be readily applied in the clinical setting by training with medical images and corresponding radiology reports that are generated routinely during clinical care, adding no additional data labeling cost. Many institutions, including ours, de-identify this data to maintain patient privacy while providing a unique setting for research on this high-quality human-annotated data. Recent ethical viewpoints describe how large-scale already-acquired clinical data could be responsibly used for secondary purposes such as training ML algorithms, to ensure that the data are used for the benefit of future patients and society Larson et al. (2020). Following such frameworks, VLMs like MedCLIP Wang et al. (2022), BiomedCLIP Zhang et al. (2023a), LLaVA-Med Li et al. (2023a), Med-Flamingo Moor et al. (2023), Med-PaLM M Tu et al. (2023), RadFM Wu et al. (2023), XrayGPT Thawkar et al. (2023), ELIXR Xu et al. (2023), RoentGen Chambon et al. (2022); Bluethgen et al. (2024), CheXagent Chen et al. (2024), and MAIRA-1 Hyland et al. (2023) demonstrate the potential of VLMs applied within the radiology domain.
@@
    We develop a training strategy and foundation model called MerlinNamed after the legendary figure said to be able to perceive everything in the present, our model Merlin simultaneously processes the entirety of 3D information in a CT volume. that leverages the structured and unstructured data within hospitals to train an abdominal CT visual model. Using this training strategy, we train a vision-language model on a single GPU using paired CTs (6,387,231 images from 15,331 CTs), EHR diagnosis codes (1,839,559 codes), and radiology reports (6,036,645 tokens). Merlin allows processing the entire 3D voxel data in a CT image at once.
@@
    We evaluate Merlin on a comprehensive set of 6 task types and 752 individual tasks, covering a variety of diagnostic, prognostic, and quality-related tasks. The non-adapted (off-the-shelf tasks) include zero-shot findings classification (31 findings), phenotype classification (692 phenotypes), and zero-shot cross-modal retrieval (image to findings and image to impressions), while model adapted tasks include 5-year disease prediction (6 diseases), radiology report generation, and 3D semantic segmentation (20 organs).
@@
    We perform internal validation on 5,137 CTs, as well as external validation on 44,098 CTs from three external sites spanning both abdomen and chest anatomies, and on two publicly available datasets focused on abdominal CT (VerSe Liebl et al. (2021) and TotalSegmentator Wasserthal et al. (2023)). We demonstrate that our singular Merlin model outperforms carefully chosen task-specific baselines on our suite of benchmarking tasks.
@@
    We compare Merlin to state-of-the-art finetuned 2D VLMs, 2D-to-3D lifted VLMs, and 3D vision-only models and show that (i) Merlin’s vision-language pretraining outperforms vision-only pretraining, (ii) finetuning improves the performance of 2D and 2D-to-3D VLMs though they underperform relative to Merlin, and (iii) Merlin outperforms other baselines in both data scarce and fully supervised settings.
  5. 5.
@@
  6. 6.
    We release our trained models, code, and dataset. Our dataset, the Merlin Abdominal CT dataset, has been manually reviewed to ensure that there is no personal health information in 25,494 CT images and corresponding radiology reports.
@@ 2 Results
We present results across 6 evaluation task types comprising 752 individual tasks, that can be performed off-the-shelf without adaptation (Figure 1b - Figure 1d), or with additional adaptation (Figure 1e - Figure 1g). The non-adapted tasks that we evaluate Merlin on include zero-shot classification (31 classification tasks; Figure 1b), phenotype classification (692 classification tasks; Figure 1c), and zero-shot cross-modal retrieval (image to findings and image to impressions; Figure 1d). The adapted tasks that we evaluate Merlin on include 5-year disease prediction (6 classification tasks; Figure 1e), radiology report generation (Figure 1f), and 3D segmentation (20 organs; Figure 1g). We report that training Merlin required approximately 160 hours on a single NVIDIA A6000 GPU.
Figure 1: Overview of Merlin training and evaluation. (a) Merlin training strategy. Diagnosis codes from the EHR are used as labels for Merlin training, with a binary cross entropy loss. Radiology reports are also used for training, with an InfoNCE loss Oord et al. (2018). Training with diagnosis codes and radiology reports is either staged or performed in a multi-task manner. Merlin is then evaluated on non-adapted tasks that can be performed without any architectural or weight modifications. These include (b) zero-shot findings classification, (c) phenotype classification, and (d) zero-shot cross-modal retrieval. Adapting Merlin enables us to perform (e) 5-year disease prediction, (f) radiology report generation, and (g) 3D semantic segmentation. (h) Illustration of Merlin’s potential clinical applications. Merlin can assist radiologists by generating radiology reports and segmenting organs from CT scans. Additionally, it can facilitate the retrieval of radiology reports from patients with similar CT scans, aid clinicians in identifying potential future disease for patients, and help minimize billing errors by assigning ICD codes.
@@ 2.1 Zero-shot Findings Classification
Zero-shot findings classification assesses Merlin’s ability to classify the presence of common imaging findings based on text prompts that a user can develop, which are likely distinct from prompts seen during training. In Figure 2b we evaluate Merlin zero-shot classification across 30 abdominal CT findings on our internal and external clinical datasets. Merlin achieves an average F1 score of 0.741 (95%CI [0.727-0.755]) on the internal validation dataset and an average F1 score of 0.647 (95%CI [0.607-0.678]) on the external validation dataset, significantly outperforming off-the-shelf 2D OpenCLIP Cherti et al. (2022) with K=1 pooling and finetuned 2D BioMedCLIP Zhang et al. (2023b) with average pooling (p < 0.001) in both settings (Figure 2b). Additional zero-shot ablation results for OpenCLIP and BiomedCLIP are presented in Supplemental Figure 10. Qualitatively, we observe in Figure 2c that Merlin’s external performance remains high on diseases with coarse-grained or salient features, e.g. pleural effusion, splenomegaly, ascites, surgically absent gallbladder, prostatomegaly, anasarca, and abdominal aortic aneurysm. Performance expectedly decreases on more challenging findings that require more subtle and fine-grained features, e.g. appendicitis, metastatic disease, lymphadenopathy, and free air. We also plot Merlin performance without radiology report splitting (W/O splitting in Figure 2c). Radiology report splitting refers to splitting the radiology reports into sections describing different anatomical structures for subsequent contrastive learning (e.g. "liver: normal" and "vasculature: patent"). Without radiology report splitting, Merlin achieves an average F1 score of 0.656 (95%CI [0.640, 0.671]), similar to the Merlin external evaluation performance. On a separate external dataset (VerSe Löffler et al. (2020) vertebral fracture detection), Merlin achieves a zero-shot F1 of 0.767 (95%CI [0.623-0.867]). In Figure 2d, we establish quantitative relationships to assess how expanding the pretraining dataset would improve zero-shot classification. This analysis helps determine the extent of pretraining data necessary for obtaining a specified zero-shot classification performance.
@@
We further reformulated the zero-shot classification task into a supervised zero-shot findings classification task to compare zero-shot Merlin against supervised Merlin baselines using 10% and 100% pretraining data (Figure 8d). In line with prior work from the original CLIP paper Radford et al. (2021a), zero-shot Merlin outperformed all supervised baselines in both data-scarce (10% training data) and fully-labeled (100% training data) supervised experiments, achieving a 45.0% increase in F1 score with 10% training data and a 29.0% increase with 100% training data (Figure 8d). Specifically, zero-shot Merlin outperformed supervised Merlin in the 100% training data setting, achieving a 16.0% increase in F1 score (0.741; 95% CI [0.727-0.755] versus 0.641; 95% CI [0.574-0.701]). Additionally, Merlin trained with 100% of the data achieved a 13.0% increase in F1 score than Merlin trained on 10% of the data, which had an F1 score of 0.577 (95% CI [0.516-0.634]).
Figure 2: Zero-shot findings classification. (a) Depicts how zero-shot classification is performed where text embeddings from disease presence prompts and disease absence prompts are compare to the image embedding. (b) We compare the performance of off-the-shelf 2D OpenCLIP Cherti et al. (2022) with K=1 pooling, finetuned 2D BioMedCLIP Zhang et al. (2023b) with average pooling, Merlin on an internal dataset, and Merlin without radiology report splitting. We further evaluate Merlin on an external clinical dataset and the VerSe Löffler et al. (2020) external fracture detection dataset. (c) Performance of finetuned 2D BioMedCLIP, Merlin, and Merlin without report splitting on the internal dataset, as well as Merlin on the external dataset, across 30 findings assessed on abdominal CT scans. (d) Merlin zero-shot classification performance improves with increasing pretraining dataset size. (e) An ablation study across various aspects of Merlin’s pretraining strategy. "Rpt." is shorthand for "report" and indicated training with radiology reports only. Staged (Stg.) refers to performing weakly supervised training with EHR in a first training stage and then training with radiology reports in a second stage. This is in contrast to multi-task learning (MTL) where EHR and radiology reports are used for training simultaneously.
@@ 2.2 Phenotype Classification
In Figure 3b we compute data scaling law curves to assess how Merlin performance improves with respect to training data. Similar to zero-shot classification, we find that increasing training data improves performance.
@@
Figure 3: Phenotype classification. (a) Average AUROC performance for the top 20 phenotype groups listed in order of prevalence (black line). (b) Data scaling law experiments that measure how average AUROC and AUPRC scale across the 692 phenotypes as the amount of pretraining data varies. (c) Average AUROC as a function of model stem hyper-parameters. We find that a smaller receptive field yields better performance. (d) Average AUROC (left chart) and AUPRC (right chart) across all 692 phenotypes, the top quartile of 173 phenotypes, and the botton quartile of 173 phenotypes across several baseline models. All baseline models are trained using the phenotypes in the pretraining dataset. The dashed lines denote random chance performance. Note that Merlin, which uses the best performing backbone of ResNet152, is further trained using radiology reports.